# Neural Guided Diffusion Bridges

### v2

@@ 6 Conclusions and limitations
We propose the neural guided diffusion bridge, a novel method for simulating diffusion bridges that enhances guided proposals through variational inference, eliminating the need for MCMC or SMC. This approach enables efficient independent sampling with comparable quality in challenging tasks where existing score-learning-based methods struggle. Extensive experiments, including both quantitative and qualitative evaluations, validate the effectiveness of our method. However, as the framework is formulated variationally and optimized by minimizing $\mathrm{D_{KL}}{(\mathbb{P}^{\bullet}_{\theta}||\mathbb{P}^{\star})}$, it exhibits mode-seeking behaviour, potentially limiting its ability to explore all modes compared to running multiple MCMC chains. Despite this limitation, our method provides a computationally efficient alternative to guided proposals, particularly in generating independent conditioned samples.
Our approach focuses on better approximating the drift of the conditioned process while keeping the guiding term that ensures the process hits $v$ at time $T$ relatively simple. In future work, one could try to jointly learn $\vartheta_{\theta}$ and $B$ using variational inference. Another venue of future research can be to extend our approach to conditioning on partial observations at multiple future times.
@@ Acknowledgements
The work presented in this paper was supported by the Villum Foundation Grant 40582, and the Novo Nordisk Foundation grant NNF18OC0052000.
@@ A.1 Preconditioned Crank-Nicolson
A more general version of using pCN to sample $X^{\star}$ is illustrated in Mider et al. [2021, Section 4], where multiples states can be observed at times $0=t_{0}<t_{1}<\dots<t_{n}$, while in our case, we suppose only the endpoint state is observed, leading to a simpler version of Mider et al. [2021, Algorithm 4.1], as Algorithm 2.
Algorithm 2 Preconditioned Crank-Nicolson
@@
2:Solve Equation (12) on $\mathcal{T}$, obtain $\{\tilde{H}(t_{m})\},\{\tilde{r}(t_{m},\cdot)\}$ using Equation (11).
@@
4:Solve Equation (6) on $\mathcal{T}$ with $w=\{w_{t_{m}}\}$, obtain $y=\{y_{t_{m}}\}$.
@@
8:     Solve Equation (6) on $\mathcal{T}$ with $z=\{z_{t_{m}}\}$, obtain $y^{\circ}=\{y^{\circ}_{t_{m}}\}$.
9:     Compute $A=\Psi(y^{\circ})/\Psi(y)$ with $\{y^{\circ}_{t_{m}}\}|$ and $\{y_{t_{m}}\}$ using Equation (10a).
@@ A.2 Proof of Theorem 2
$\mathrm{D}_{\text{KL}}{(\mathbb{P}^{\bullet}_{\theta}||\mathbb{P}^{\star})}$ $=\mathbb{E}^{\bullet}\left[\log\left(\frac{\mathrm{d}\mathbb{L}^{\bullet}_{\theta}}{\mathrm{d}\mathbb{L}^{\star}}\right)(X)\right]$
$=\mathbb{E}^{\bullet}\left[\log\left(\frac{\mathrm{d}\mathbb{L}^{\bullet}_{\theta}}{\mathrm{d}\mathbb{L}^{\circ}}\cdot\frac{\mathrm{d}\mathbb{L}^{\circ}}{\mathrm{d}\mathbb{L}^{\star}}\right)(X)\right]$
@@
where the stochastic integral vanishes because of the martingale property of the Itô integral. The first equality follows from Equation (14). By Equation (7)
@@
Substituting Equation (33b) and Equation (34) into Equation (32a) gives
$\mathrm{D}_{\text{KL}}{(\mathbb{P}^{\bullet}_{\theta}||\mathbb{P}^{\star})}=\mathbb{E}^{\bullet}\int_{0}^{T}\left\{\frac{1}{2}\|\vartheta_{\theta}(t,X_{t})\|^{2}-G(t,X_{t})\right\}\mathrm{d}t+\log\frac{\tilde{h}(0,x_{0})}{h(0,x_{0})}=L(\theta)+\log\frac{\tilde{h}(0,x_{0})}{h(0,x_{0})}\geq 0,$ (35)
as $L(\theta)$ defined as Equation (18). ∎
@@ A.3 SDE gradients
We now derive the gradient of Equation (21) with respect to $\theta$, on a fixed Wiener realization $w^{\bullet(n)}=\{w^{\bullet(n)}_{t_{m}}\}$. As discussed, $x^{\bullet(n)}_{t_{m}}=\phi_{\theta}(w^{\bullet(n)}_{t_{m}})$ is implemented as a numerical SDE solver $f_{\theta}(w^{\bullet(n)}_{t_{m}},t_{m-1},x^{\bullet(n)}_{t_{m-1}}),m\geq 1$ that takes the previous step $(t_{m-1},x^{\bullet(n)}_{t_{m-1}})$ as additional arguments. As $x^{\bullet(n)}_{t_{m-1}}$ also depends on $\theta$, the gradient with respect to $\theta$ needs to be computed recursively. Specifically, with $x^{\bullet(n)}_{t_{m}}=f_{\theta,m}=f_{\theta}(w^{\bullet(n)}_{t_{m}},t_{m-1},x^{\bullet(n)}_{t_{m-1}})$
@@
The automatic differentiation can save all the intermediate Equation (40a) and Equation (40b), which enables to compute $\nabla_{\theta}L(\theta)$.
@@ B.2 Linear processes
Brownian bridges: When no noise is in observations, $h(0,x_{0})=\mathcal{N}(y;x-\gamma T,\sigma^{2}T)$, $\tilde{h}(0,x_{0})=\mathcal{N}(y;x,\sigma^{2}T)$, therefore, the lower bound of $L(\theta)$, $-\log\frac{\tilde{h}(0,x_{0})}{h(0,x_{0})}=(\|y-x_{0}-\gamma T\|^{2}-\|y-x_{0}\|^{2})/(2\sigma^{2}T)$. In practice, we set $\epsilon=10^{-10}$, $\vartheta_{\theta}(t,x)$ is modelled as a fully-connected neural network with 3 hidden layers and 20 hidden dimensions for each layer. The model is trained with 60,000 independently sampled full trajectories of $X^{\bullet}_{t}$. The batch size $N=20$, time step size $\delta t=0.005$, therefore in total $M=200$ time steps. The network is trained with Adam [Kingma and Ba, 2017] optimizer with an initial learning rate of $1.0e^{-3}$ and a cosine decay scheduler, and a linear learning rate warming up in the first 0.1 ratio of training iterations.
Figure 9: Results of Brownian bridges under different $\gamma$. Top row: the training loss log, the red dashed line represents the lower bound of $L(\theta)$.
@@
Ornstein-Uhlenbeck bridge: When conditioning Equation (24) on $\{LX_{T}=v\}$, the conditioned process $X^{\star}$ reads:
$\mathrm{d}X^{\star}_{t}=-\gamma\left(\frac{v}{\operatorname{sinh}(-\gamma(T-t))}-\frac{x}{\operatorname{tanh}(-\gamma(T-t))}\right)\mathrm{d}t+\sigma\mathrm{d}W_{t},$ (41)
which is obtained from the fact that
$p(T,y\mid t,x)=\sqrt{\frac{\gamma}{\pi\sigma^{2}(1-\exp(-2\gamma(T-t))}}\exp\left\{-\frac{\gamma}{\sigma^{2}}\cdot\frac{(y-x\exp(-\gamma(T-t))^{2}}{1-\exp(-2\gamma(T-t))}\right\}.$ (42)
Therefore, one can immediately see the optimal value of $\vartheta_{\theta}(t,x)$ to be
@@
$\vartheta_{\theta_{\text{opt}}}(t,x)=\frac{1}{\sigma}\left\{-\gamma\left[\frac{v}{\sinh(-\gamma(T-t))}-\frac{x}{\tanh(-\gamma(T-t))}\right]-\frac{v-x}{T-t}+\gamma\cdot x\right\},$ (43)
and the lower bound of $L(\theta)$ to be:
$-\log\frac{\tilde{h}(0,x_{0})}{h(0,x_{0})}=\frac{1}{2}\left\{\log\left(\frac{\gamma}{\pi\sigma^{2}(1-e^{-2\gamma T})}\right)-\frac{\gamma(y-x_{0}e^{-\gamma T})^{2}}{\sigma^{2}(1-e^{-2\gamma T})}-\log\left(\frac{1}{2\pi\sigma^{2}T}\right)+\frac{(y-x_{0})^{2}}{\sigma^{2}T}\right\}.$ (44)
These analytical results enable a quantitative evaluation of the performance of learned $X^{\bullet}$. The training setups are duplicated from the previous Brownian bridges, except for a deeper neural network architecture with 4 hidden layers and 20 hidden units in each layer.
Figure 10: Results of Ornstein-Uhlenbeck bridges under different $\gamma$. Top row: the training loss log, the red dashed line represents the lower bound of $L(\theta)$.
@@ B.3 Cell diffusion process
For the benchmark tests, we adapt the published implementations of the corresponding methods to fit into our test framework with possibly minor modifications. Specifically, the original guided proposal codebase is implemented with Julia in  [^1] https://juliapackages.com/p/bridge, we rewrite it in JAX [Bradbury et al., 2018]; the score matching bridge repository is published in  [^2] https://github.com/jeremyhengjm/DiffusionBridge, which is written with PyTorch [Paszke et al., 2019], we also adapt it into JAX for convenience, additionally, as also reported by the authors, Heng et al. [2022] introduces two score-matching-based bridge simulation schemes, reversed and forward simulation, and the forward simulation relies on the reversed simulation, and learning from approximated reversed bridge can magnifies the errors due to progressive accumulation. Therefore, we only compare our method with the reversed bridge learning to avoid error accumulations; the adjoint bridge is implemented in  [^3] https://github.com/libbylbaker/forward_bridge, already in JAX, so we directly use the existing implementation without modification.
Normal event: We still consider a no-observation-noise case by setting $\epsilon=10^{-10}$, $\vartheta_{\theta}$ is modelled as a fully-connected network with 4 hidden layers and 32 hidden dimensions per layer, activated by tanh. We train the model with 100,000 independent samples of $X^{\bullet}$ by a Adam optimizor with initial learning rate of $5e^{-4}$, batch size $N=20$, time step $\delta t=0.01$, therefore, in total $M=400$ time steps. For the guided proposal, we set $\eta=0.98$, and run one chain of MCMC for 10,000 updates, obtaining $22.69\%$ acceptance rate. The burn-in period is set as the first 4,000 steps. After burn-in, we subsample the updating history by taking every 200-th sample, therefore selecting out 30 samples. In the following experiments, if not explicitly stated, all the samples from the guided proposal are similarly obtained from one chain of MCMC by taking proper thinning ratios. For the score matching, we find the original neural network architecture in the repository performs poorly in our case, so we refer to the architecture used in Baker et al. [2024a], with 4 hidden layers and 32 hidden dimensions per layer, combining with the sinusoidal embedding [Vaswani et al., 2017] for the temporal dependency. For the adjoint bridge, we directly deploy the original architecture used in Baker et al. [2024b]. To obtain the reference dynamics, we generate 100,000 independent samples of $X$, then filter these by the condition $\|LX_{T}-v\|\leq 0.01$, obtaining 151 valid samples, and only the first 30 samples are shown in grey.
@@
Rare event: The setups for conditioning on rare events of the neural guided bridge are replicated from the previous normal event case, except for the MCMC is running with sightly increased tuning parameter $\eta=0.99$ and obtains an acceptance rate of $24.50\%$. For the score matching bridge, since we only use the reversed bridge, whose learning is independent of the endpoints, we directly deploy the trained score approximation from the previous case; For the adjoint process, we fix the neural network architecture and training scheme, changing only the conditions. To obtain the reference, we repeat the same procedure of sampling 100,000 independent trajectories and filtering by the condition, however, we find none of the samples matches the condition we put, which reflects that $v=[1.0,-0.1]^{T}$ is indeed a probably rare event and becomes challenging.
Multi-modality: All the neural network architectures and training schemes are the same as in previous examples. $\delta t=0.01$ and $M=500$ therefore. For the guided proposal, we run one chain for 10,000 iterations with $\eta=0.9$, obtaining $24.93\%$ acceptance rate. Similarly to the rare event case, no valid samples can be found through 100,000 brute-forced samples of the original process. Figure 11 shows marginal distributions of the original and conditioned processes, where more than one peak of marginal distribution density can be observed at time $t=3.0,4.0$. Neither the neural guided process nor the guided proposal can recover all three modes which are indicated by the original process. However, one can always run multiple independent chains to cover all the modes, as suggested in Figure 12. Compared to the guided proposal, even though the neural bridge can not recover all the modes, it trades with fast sample speed. Also, additional MCMC updates can be further executed on trained neural bridges to obtain high-quality samples.
Figure 11: Comparison of marginal distributions at different time slices of the original process, guided proposal and learned neural bridge. The histograms are made from 1,000 samples individually either drawn independently from the neural guided bridge and original process or subsampled from the guided proposal updating history. Top row: the first component of $X$; Bottom row: the second component of $X$.
Figure 12: Marginal distributions of two independent MCMC chains, where they compensate each other and cover all three modes.

### v3

@@ 6 Conclusions and Limitations
We propose the neural guided diffusion bridge, a novel method for simulating diffusion bridges that enhances guided proposals through variational inference, eliminating the need for MCMC or SMC. This approach enables efficient independent sampling with comparable quality in challenging tasks where existing score-learning-based methods struggle. Extensive experiments, including both quantitative and qualitative evaluations, validate the effectiveness of our method. However, as the framework is formulated variationally and optimized by minimizing $\mathrm{D_{KL}}{(\mathbb{P}^{\bullet}_{\theta}||\mathbb{P}^{\star})}$, it exhibits mode-seeking behaviour, potentially limiting its ability to explore all modes compared to running multiple MCMC chains. Despite this limitation, our method provides a computationally efficient alternative to guided proposals, particularly in generating independent samples from the conditioned process.
Our approach focuses on better approximating the drift of the conditioned process while keeping the guiding term that ensures the process hits $v$ at time $T$ relatively simple. In future work, an interesting direction to obtain improved results consists in trying to jointly learn $\vartheta_{\theta}$ and the parameters of the linear process ($\tilde{\sigma}$, $B$, $\beta$). Also, as the gradient updating relies on backpropagating through the whole numerical SDE solvers, techniques such as the stochastic adjoint sensitivity method [Li et al., 2020] and adjoint matching [Domingo-Enrich et al., 2025] can be introduced to improve the efficiency of the computation. Another venue of future research consists of extending our approach to conditioning on partial observations at multiple future times.
@@ Acknowledgements
We thank Thorben Pieper-Sethmacher for his valuable feedbacks on the paper. The work presented in this paper was supported by the Villum Foundation Grant 40582, and the Novo Nordisk Foundation grant NNF18OC0052000.
@@ A.1 Preconditioned Crank-Nicolson
Algorithm 2 is a special case of Algorithm 4.1 of Mider et al. [2021].
Algorithm 2 Preconditioned Crank-Nicolson scheme for guided proposals
@@
2:  Solve Equation 15 on $\mathcal{T}$, obtain $\{\tilde{H}(t_{m})\},\{\tilde{r}(t_{m},\cdot)\}$ using Equations 13 and 14.
@@
4:  Solve Equation 7 on $\mathcal{T}$ with $w=\{w_{t_{m}}\}$, obtain $y=\{y_{t_{m}}\}$.
@@
8:     Solve Equation 7 on $\mathcal{T}$ with $z=\{z_{t_{m}}\}$, obtain $y^{\circ}=\{y^{\circ}_{t_{m}}\}$.
9:     Compute $A=\Psi(y^{\circ})/\Psi(y)$ with $\{y^{\circ}_{t_{m}}\}$ and $\{y_{t_{m}}\}$ using Equation 11.
@@ A.2 Proof of Theorem 4
$\mathrm{D}_{\text{KL}}{(\mathbb{P}^{\bullet}_{\theta}||\mathbb{P}^{\star})}$ $=\mathbb{E}^{\bullet}\left[\log\left(\frac{\mathrm{d}\mathbb{L}^{\bullet}_{\theta}}{\mathrm{d}\mathbb{L}^{\star}}\right)(X)\right]=\mathbb{E}^{\bullet}\left[\log\left(\frac{\mathrm{d}\mathbb{L}^{\bullet}_{\theta}}{\mathrm{d}\mathbb{L}^{\circ}}\cdot\frac{\mathrm{d}\mathbb{L}^{\circ}}{\mathrm{d}\mathbb{L}^{\star}}\right)(X)\right]$
@@
where the stochastic integral vanishes because of the martingale property of the Itô integral. The first equality follows from Equation 18. By Equation 8
@@
Substituting Equation 35b and Equation 36 into Equation 34a gives
$\mathrm{D}_{\text{KL}}{(\mathbb{P}^{\bullet}_{\theta}||\mathbb{P}^{\star})}=\mathbb{E}^{\bullet}\int_{0}^{T}\left\{\frac{1}{2}\|\vartheta_{\theta}(t,X_{t})\|^{2}-G(t,X_{t})\right\}\mathrm{d}t-\log\frac{\tilde{h}(0,x_{0})}{h(0,x_{0})}=L(\theta)-\log\frac{\tilde{h}(0,x_{0})}{h(0,x_{0})}\geq 0,$ (37)
with $L(\theta)$ as defined in Equation 20. ∎
@@ A.3 SDE Gradients
We now derive the gradient of Equation 23 with respect to $\theta$, on a fixed Wiener realization $w^{\bullet(n)}=\{w^{\bullet(n)}_{t_{m}}\}$. As discussed, $x^{\bullet(n)}_{t_{m}}=\phi_{\theta}(w^{\bullet(n)}_{t_{m}})$ is implemented as a numerical SDE solver $f_{\theta}(w^{\bullet(n)}_{t_{m}},t_{m-1},x^{\bullet(n)}_{t_{m-1}}),m\geq 1$ that takes the previous step $(t_{m-1},x^{\bullet(n)}_{t_{m-1}})$ as additional arguments. As $x^{\bullet(n)}_{t_{m-1}}$ also depends on $\theta$, the gradient with respect to $\theta$ needs to be computed recursively. Specifically, with $x^{\bullet(n)}_{t_{m}}=f_{\theta,m}=f_{\theta}(w^{\bullet(n)}_{t_{m}},t_{m-1},x^{\bullet(n)}_{t_{m-1}})$
@@
The automatic differentiation can save all the intermediate Equation 42a and Equation 42b, which enables to compute $\nabla_{\theta}L(\theta)$.
@@ B.2 Linear Processes
Brownian bridges: If $\mathrm{d}X_{t}=\gamma\mathrm{d}t+\sigma\mathrm{d}W_{t}$, then
$\log h(t,x)=\log p(T,v\mid t,x)=-\frac{1}{2}\log(2\pi\sigma^{2}(T-t))-\frac{(v-x-\gamma(T-t))^{2}}{2\sigma^{2}(T-t)}.$ (43)
If $\mathrm{d}\tilde{X}_{t}=\sigma\mathrm{d}W_{t}$, then $\log\tilde{h}(t,x)$ is obtained by taking $\gamma=0$ in the preceding display.
Therefore, in this case we can compute the lower bound on the loss: $L(\theta)\geq\log\frac{\tilde{h}(0,x_{0})}{h(0,x_{0})}=\frac{(v-x-\gamma T)^{2}-(v-x)^{2}}{2\sigma^{2}T}$. Moreover, the optimal map $\vartheta_{\theta_{\text{opt}}}$ is given by $\vartheta_{\theta_{\text{opt}}}(t,x)=\sigma(\partial_{x}\log h(t,x)-\partial_{x}\log\tilde{h}(t,x))=-\frac{\gamma}{\sigma}$.
@@
In the numerical experiment, we took $\epsilon=10^{-5}$. The map $\vartheta_{\theta}$ is modeled by a fully connected neural network with 3 hidden layers and 20 hidden dimensions for each layer. The model is trained with 25,000 independently sampled full trajectories of $X^{\bullet}$. The batch size was taken to be $N=50$ and the time step size $\delta t=0.002$, leading to in total $M=500$ time steps. The network was trained using the Adam [Kingma and Ba, 2017] optimizer with learning rate $0.001$.
Figure 2: Evaluations of trained neural networks $\vartheta_{\theta}(t,x)$ against the optimal maps $\vartheta_{\theta_{\text{opt}}}(t,x)$ under different settings of $\gamma,\sigma$, where the background colour intensities indicate the absolute error $|\vartheta_{\theta}-\vartheta_{\theta_{\text{opt}}}|$. In each panel, 10 independent samples from the guided proposal are shown in grey to indicate the sampling regions where one expects the error to be smallest.
Figure 3: Comparison of marginal distributions at different time slices of the learned neural bridge (n.b.) and analytical true Brownian bridge (t.b.) under the setting $\gamma=1.0,\sigma=1.0$, conditioned on $v=0.0$. The purple bars show the absolute error (abs. error) of each bin between the learned neural bridge and the histograms obtained by forward sampling from the true bridge process, which are expected to be low when two distributions are close in terms of their shapes. The histograms are made from 50,000 independent trajectory samples.
@@
Ornstein-Uhlenbeck bridge: When conditioning Equation 26 on $v$, the conditioned process $X^{\star}$ satisfies the SDE
$\mathrm{d}X^{\star}_{t}=\left\{\gamma(\mu-X^{\star}_{t})+\frac{2\gamma e^{-\gamma(T-t)}}{1-e^{-2\gamma(T-t)}}\left[(v-\mu)-e^{-\gamma(T-t)}(X^{\star}_{t}-\mu)\right]\right\}\mathrm{d}t+\sigma\mathrm{d}W_{t},$ (44)
which is obtained from the transition density given by Equation 26 is:
$p(T,y\mid t,x)=\frac{1}{\sqrt{2\pi\Sigma^{2}_{t,T}}}\exp\left(-\frac{(v-m_{t,T}(x))^{2}}{2\Sigma^{2}_{t,T}}\right),$ (45a)
$m_{s,t}(x)=\mu+(x-\mu)e^{-\gamma(t-s)},$ (45b)
$\Sigma^{2}_{s,t}=\frac{\sigma^{2}}{2\gamma}\left(1-e^{-2\gamma(t-s)}\right).$ (45c)
Since the auxiliary process is chosen the same as the Brownian case, one can easily show the optimal value of $\vartheta_{\theta}(t,x)$ to be
@@
$\vartheta_{\theta_{\text{opt}}}(t,x)=\frac{2\gamma e^{-\gamma(T-t)}}{\sigma(1-e^{-2\gamma(T-t)})}\left[(v-\mu)-e^{-\gamma(T-t)}(x-\mu)\right]+\frac{(v-x)}{\sigma(T-t)}.$ (46)
Therefore, the lower bound on $\theta\mapsto L(\theta)$ is given by
$\log\frac{\tilde{h}(0,x_{0})}{h(0,x_{0})}=-\frac{1}{2}\log(2\pi\sigma^{2}(T-t))-\frac{(v-x)^{2}}{2\sigma^{2}(T-t)}+\frac{1}{2}\log(2\pi\Sigma^{2}_{t,T})+\frac{(v-m_{t,T}(x))^{2}}{2\Sigma^{2}_{t,T}}.$ (47)
We repeated the same numerical and experimental settings as the previous Brownian example, except for training with 50,000 samples to obtain better results.
Figure 4: Evaluations of trained neural networks $\vartheta_{\theta}(t,x)$ against the optimal maps $\vartheta_{\theta_{\text{opt}}}(t,x)$ under different settings of $\gamma,\mu,\sigma$, where the background colour intensities indicate the absolute error $|\vartheta_{\theta}-\vartheta_{\theta_{\text{opt}}}|$. In each panel, 10 independent samples from the guided proposal are shown in grey to indicate the sampling regions where one expects the error to be smallest. Except the rightmost setting with $\gamma=1.7,\mu=0.0,\sigma=0.3$, all the processes are conditioned on $v=1.0$, whereas in the rightmost case, the process is conditioned on $v=0.0$.
Figure 5: Comparison of marginal distributions at different time slices of the learned neural bridge (n.b.) and analytical true OU bridge (t.b.) under the setting $\gamma=1.7,\mu=1.0,\sigma=0.3$, conditioned on $v=1.0$. The purple bars show the absolute error (abs. error) of each bin. The histograms are made from 50,000 independent trajectory samples.
(a) Brownian bridges
(b) OU bridges
Figure 6: Loss curves of training Brownian and OU bridges. For Brownian bridges, the SDE parameters are taken as $\gamma=1.0,\sigma=1.0$, for the OU bridges, the parameters are taken as $\gamma=1.7,\mu=1.0,\sigma=0.3$ unless any of them is specified.
@@ B.3 Cell Diffusion Process
For the benchmark tests, we adapt the published guided proposal implementations of the corresponding methods to fit into our test framework with possibly minor modifications. Specifically, the original guided proposal codebase is implemented with Julia in  [^1] https://juliapackages.com/p/bridge, we rewrite it in JAX [Bradbury et al., 2018]; the score matching bridge repository is published in  [^2] https://github.com/jeremyhengjm/DiffusionBridge. additionally, as also reported by the authors, Heng et al. [2022] introduces two score-matching-based bridge simulation schemes, reversed and forward simulation, and the forward simulation relies on the reversed simulation, and learning from approximated reversed bridge can magnifies the errors due to progressive accumulation. Therefore, we only compare our method with the reversed bridge earning to avoid error accumulations; the adjoint bridge is implemented in  [^3] https://github.com/libbylbaker/forward_bridge.
Normal event: We took $\epsilon^{2}=10^{-10}$ and $\vartheta_{\theta}$ is modeled as a fully-connected network with $3$ hidden layers and 32 hidden dimensions per layer, activated by LipSwish. We trained the model for $5,000$ gradient descent updating iterations. In each step, we sampled a batch of $100$ independent trajectories of $X^{\bullet}$ under the current $\theta$. The numerical sampling time step size is $\delta t=0.01$. Therefore, in total $M=400$ discrete steps of a single trajectory. The Adam optimizer with learning rate of $31e-$ 3 was used for optimization. For the guided proposal, we set $\eta=0.98$, and ran one MCMC chain for 10,000 iterations. This resulted in an acceptance percentage of $21.38\%$. The initial $5,000$ iterations were considered as burn-in samples. After burn-in, we collected the samples and subsampled them by taking every 133 samples to obtain 30 samples. In the following experiments, if not explicitly stated, all the samples from the guided proposal are similarly obtained from one MCMC chain by subsampling from the outputs. For the score matching and adjoint forward methods, we deployed the given network structures provided in their codebases As a reference, we sampled from the forward process until we obtained $30$ samples satisfying the inequality $\|LX_{T}-v\|\leq 0.01$. These samples are shown in grey.
(a) neural bridge
(b) guided proposal
(c) score matching
(d) adjoint forward
Figure 7: Visualization of 30 simulated bridge trajectories under normal event conditioning, using various sampling methods. All samples are independently drawn, except those generated by the guided proposal. As a reference, in grey, we added trajectories from the unconditional forward process, with samples that satisfy the condition $\|LX_{T}-v\|>0.01$ rejected.
@@
Rare event: The setups for conditioning on rare events of the neural guided bridge are replicated from the previous normal event case, except for the MCMC is running with sightly increased tuning parameter $\eta=0.99$ and for 20,000 iterations and the first 10,000 iterations is discarded as the burn-in period. The acceptance percentage is $22.29\%$. For the score matching, since we only use the reversed bridge, where learning is independent of the event we condition on , we directly deploy the trained score approximation from the previous case. For the adjoint forward, we fix the neural network architecture and training scheme, changing only the conditioned target.
As a reference, we sampled the forward process $100,000$ times. This time however, none of the samples ended near the point we condition on, confirming we are dealing with a rare event.
(a) neural bridge
(b) guided proposal
(c) score matching
(d) adjoint forward
Figure 8: Visualization of 30 simulated bridge trajectories under rare event conditioning, using various sampling methods. All samples are independently drawn, except those generated by the guided proposal. Contrary to Figure 7, no reference trajectories were added as extensive forward simulation of the process did not yield any samples satisfying $\|LX_{T}-v\|<0.01$.
Multi-modality: All neural network architectures and training settings match those used in previous examples. We set $\delta t=0.01$ and $M=500$. For the guided proposal, a single chain is run for $50,000$ iterations with $\eta=0.9$, discarding the first $20,000$ as burn-in. The acceptance percentage is $26.81\%$. As in the rare event case, no valid samples were found from forward simulating $100,000$ times the (unconditioned) forward process.
Figure 10 shows marginal distributions for both unconditioned and conditioned processes, using the guided proposal and the neural bridge. At $t=3.0$ and $t=4.0$, multiple peaks appear in the marginal densities, suggesting multi-modality. This is consistent with the unconditioned sampling, where multiple modes are also visible. However, neither score matching nor the adjoint forward method captures these modes. In contrast, the neural bridge and guided proposal yield similar marginals that accurately reflect the multi-modal nature of the process.
(a) neural bridge
(b) guided proposal
(c) score matching
(d) adjoint forward
Figure 9: Visualization of 30 simulated bridge trajectories under multi-modal event conditioning, using various sampling methods. All samples are independently drawn, except those corresponding to the guided proposal.
Figure 10: Marginal distributions of $X$ at selected time points for three settings: the unconditioned process (unc.), the neural bridge (n.b.), and the guided proposal sampled with pCN (g.p.). All trajectories start at $x_{0}=\begin{bmatrix}-1&-1\end{bmatrix}^{\top}$ and, The neural bridge and the guided proposal are conditioned to reach $v=\begin{bmatrix}-1&-1\end{bmatrix}^{\top}$ at the terminal time $T=5$. Each histogram is based on $10,000$ samples—independent draws for unc. and n.b. and subsampled MCMC draws for g.p. The first and second components of X are displayed in the top and bottom rows, respectively.